### v1

@@ Appendix F Interpretability
Region 1 Region 2 Corrected p-value Disorder/HC Cohen’s d
Dorsal Area 24d Area PGs $<0.001$ ASD 6.950
Area p32 Area PGs $<0.001$ ASD 6.565
Ventral Area 6 Area 8 Anterior ventral $<0.001$ ASD 6.479
Area p32 Area 8 Anterior ventral $<0.001$ ASD 6.313
Area 9 Middle Area 8 Anterior ventral $<0.001$ ASD 6.135
Area 9-46 dorsal Area PGs $<0.001$ ASD 6.057
Area ventral 23 a+b Area 8 Anterior ventral $<0.001$ ASD 6.041
Supplementary and Cingulate Eye Field Area PGs $<0.001$ ASD 5.959
Area 31pd Area 8 Anterior ventral $<0.001$ ASD 5.899
ParaHippocampal Area 3 Area 8Ad $<0.001$ ASD 5.819
Area 31 anterior Area PGs $<0.001$ ASD 5.671
Orbital Frontal Complex Area 47 medial $<0.001$ ASD 5.531
Area a24 Area 8 Anterior ventral $<0.001$ ASD 5.509
PreSubiculum Area PGs $<0.001$ ASD 5.178
Area 31p ventral Area 8 Anterior ventral $<0.001$ ASD 5.171
Area 23 dorsal Area PGs $<0.001$ ASD 5.107
Area V3A Area 47 medial $<0.001$ ASD 4.942
Area STSd anterior Area PGs $<0.001$ ASD 4.881
Area TemporoParietoOccipital Junction 3 Area 8Ad $<0.001$ ASD 4.745
Area 46 Area 47 medial $0.001$ ASD 4.683
Area IFSp Area 47 medial $0.001$ ASD 4.669
Area 10 rostral Area 8 Anterior ventral $0.001$ ASD 4.589
Area posterior 24 Area TE1 Middle $0.001$ ASD 4.574
Area posterior 24 Area 8 Anterior ventral $0.001$ ASD 4.550
Primary Visual Cortex Area 47 medial $0.001$ ASD 4.533
Area Posterior 24 prime Area 47 medial $0.001$ ASD 4.512
Premotor Eye Field Area 47 medial $0.001$ ASD 4.507
Ventral Area 6 Area PGs $0.001$ ASD 4.502
Inferior 6-8 Transitional Area Area 47 medial $0.001$ ASD 4.456
PreSubiculum Area 8 Anterior ventral $<0.001$ ASD 4.430
Area 9-46 dorsal Area 8 Anterior ventral $<0.001$ ASD 4.377
VentroMedial Visual Area 1 Area 47 medial $0.001$ ASD 4.365
Area 33 prime Area 47 medial $0.001$ ASD 4.293
Second Visual Area Area 47 medial $0.001$ ASD 4.274
Para-Insular Area Area 8 Anterior ventral $0.001$ ASD 4.227
ParaHippocampal Area 3 Area 8 Anterior ventral $0.001$ ASD 4.222
Area PFm Complex Area 8 Anterior ventral $0.001$ ASD 4.222
Third Visual Area Area 47 medial $0.001$ ASD 4.206
Area Lateral Occipital 3 Area 47 medial $0.001$ ASD 4.196
Area PGi Area 8 Anterior ventral $0.001$ ASD 4.193
Area p32 prime Area PGs $0.001$ ASD 4.178
Area 33 prime Area 8 Anterior ventral $0.001$ ASD 4.176
Superior Frontal Language Area Area 8 Anterior ventral $<0.001$ ASD 4.176
Supplementary and Cingulate Eye Field Area 8 Anterior ventral $0.001$ ASD 4.166
Middle Insular Area Area PGs $0.001$ ASD 4.161
Area dorsal 32 Area 8 Anterior ventral $<0.001$ ASD 4.079
Area anterior 32 prime Area 8 Anterior ventral $0.001$ ASD 4.073
Area posterior 24 Area 47 medial $0.001$ ASD 4.065
Area Lateral IntraParietal ventral Area 47 medial $0.002$ ASD 4.046
Area IFJp Area 47 medial $0.001$ ASD 4.034
Anterior 24 prime Area 47 medial $0.002$ ASD 4.026
Inferior 6-8 Transitional Area Area TE1 Middle $0.002$ ASD 4.011
Area 11 lateral Area 47 medial $0.002$ ASD 3.987
ParaHippocampal Area 3 Area 47l (47 lateral) $0.001$ ASD 3.980
Area IntraParietal 1 Area 8 Anterior ventral $0.002$ ASD 3.980
Superior 6-8 Transitional Area Area TE1 Middle $0.001$ ASD 3.968
Area PH Area 47 medial $<0.001$ ASD 3.963
Area Lateral Occipital 2 Area 8 Anterior ventral $0.001$ ASD 3.956
Eighth Visual Area Inferior 6-8 Transitional Area $0.002$ ASD 3.952
Area 6 anterior Area 47 medial $0.002$ ASD 3.948
Para-Insular Area Area 47 medial $0.002$ ASD 3.918
Area s32 Area PGs $0.002$ ASD 3.913
Area OP1-SII Inferior 6-8 Transitional Area $0.002$ ASD 3.897
Area a24 Area PGs $0.002$ ASD 3.895
Area TE1 Middle Area 8 Anterior ventral $0.001$ ASD 3.895
Area Lateral Occipital 2 Area 47 medial $0.002$ ASD 3.875
Area 6 medial anterior Area TE1 Middle $0.002$ ASD 3.866
Area dorsal 23 a+b Area 8 Anterior ventral $0.002$ ASD 3.858
Supplementary and Cingulate Eye Field Area TE1 Middle $0.002$ ASD 3.843
Area TemporoParietoOccipital Junction 2 Area 8 Anterior ventral $0.002$ ASD 3.840
Area posterior 9-46 ventral Area 47 medial $<0.001$ ASD 3.825
Pirform Cortex Area 47 medial $0.002$ ASD 3.804
Area Posterior Insular 1 Area 47 medial $0.002$ ASD 3.800
Fourth Visual Area Area 47 medial $0.002$ ASD 3.798
Area a24 Dorsal area 6 $0.002$ ASD 3.789
Area anterior 10p Area TE1 Middle $0.002$ ASD 3.788
Anterior 24 prime Area TE1 Middle $0.002$ ASD 3.785
Visual area 3-B Area 47 medial $0.002$ ASD 3.757
Area s32 Dorsal area 6 $0.002$ ASD 3.752
Area TemporoParietoOccipital Junction 3 Area 8 Anterior ventral $0.002$ ASD 3.744
RetroInsular Cortex Area 8 Anterior ventral $0.002$ ASD 3.736
Area 46 Area 8 Anterior ventral $<0.001$ ASD 3.731
Perirhinal Ectorhinal Cortex Area 47 medial $0.002$ ASD 3.731
Area p32 prime Area 47 medial $0.002$ ASD 3.720
Lateral Belt Complex Area 47 medial $0.002$ ASD 3.702
Anterior 24 prime Area 8 Anterior ventral $0.002$ ASD 3.698
Area OP1-SII Primary Motor Cortex $0.002$ ASD 3.694
Area 6 medial anterior Superior 6-8 Transitional Area $0.002$ ASD 3.693
Area Lateral IntraParietal dorsal Medial Area 7P (posterior) $0.002$ ASD 3.693
Area p32 prime Area 8 Anterior ventral $0.002$ ASD 3.682
Premotor Eye Field Area 8 Anterior ventral $0.002$ ASD 3.678
Entorhinal Cortex Area 47 medial $0.003$ ASD 3.654
Area 31 anterior Area 8 Anterior ventral $0.002$ ASD 3.629
Area 7m Area 8 Anterior ventral $0.002$ ASD 3.625
Area dorsal 32 Area TG dorsal $0.003$ ASD 3.608
Area Posterior 24 prime Area 8 Anterior ventral $0.002$ ASD 3.602
Area IntraParietal 1 Medial Area 7P (posterior) $0.002$ ASD 3.593
Area dorsal 23 a+b Inferior 6-8 Transitional Area $0.003$ ASD 3.579
Ventral Area 6 Area 47 medial $0.003$ ASD 3.574
Area 31pd Area PGs $0.001$ ASD 3.556
@@
Area PHT Area posterior 9-46 ventral $0.028$ ASD 2.155
ParaBelt Complex Area posterior 9-46 ventral $0.029$ ASD 2.154
Auditory 5 Complex Area posterior 9-46 ventral $0.030$ ASD 2.123
Area 9 Posterior Area posterior 9-46 ventral $0.030$ ASD 2.122
Primary Sensory Cortex Area posterior 9-46 ventral $0.031$ ASD 2.116
Area PH Area posterior 9-46 ventral $0.031$ ASD 2.115
Area 3a Area posterior 9-46 ventral $0.031$ ASD 2.109
Area 6mp Area posterior 9-46 ventral $0.032$ ASD 2.099
Lateral Belt Complex Area posterior 9-46 ventral $0.033$ ASD 2.078
Area TA2 Area posterior 9-46 ventral $0.034$ ASD 2.065
Area TE2 posterior Area posterior 9-46 ventral $0.034$ ASD 2.057
Area V3A Area posterior 9-46 ventral $0.034$ ASD 2.057
Area PFm Complex Area posterior 9-46 ventral $0.034$ ASD 2.053
Area Lateral Occipital 2 Area posterior 9-46 ventral $0.035$ ASD 2.046
Dorsal area 6 Area posterior 9-46 ventral $0.035$ ASD 2.045
VentroMedial Visual Area 1 Area posterior 9-46 ventral $0.035$ ASD 2.039
ParaHippocampal Area 3 Area posterior 9-46 ventral $0.036$ ASD 2.038
Area IFSp Area posterior 9-46 ventral $0.036$ ASD 2.037
Primary Auditory Cortex Area posterior 9-46 ventral $0.036$ ASD 2.032
Ventral Visual Complex Area posterior 9-46 ventral $0.037$ ASD 2.023
Area 1 Area posterior 9-46 ventral $0.036$ ASD 2.018
Area PGs Area posterior 9-46 ventral $0.037$ ASD 2.009
Area IntraParietal 2 Area posterior 9-46 ventral $0.037$ ASD 2.007
Area TemporoParietoOccipital Junction 1 Area posterior 9-46 ventral $0.037$ ASD 2.004
Primary Motor Cortex Area posterior 9-46 ventral $0.038$ ASD 2.001
ParaHippocampal Area 1 Area posterior 9-46 ventral $0.039$ ASD 1.999
Auditory 4 Complex Area posterior 9-46 ventral $0.039$ ASD 1.993
Area 31a Area posterior 9-46 ventral $0.039$ ASD 1.987
Area TE1 posterior Area posterior 9-46 ventral $0.039$ ASD 1.985
Superior Temporal Visual Area Area posterior 9-46 ventral $0.039$ ASD 1.980
Area 25 Area posterior 9-46 ventral $0.040$ ASD 1.978
Area V6A Area posterior 9-46 ventral $0.039$ ASD 1.977
Area OP2-3-VS Area posterior 9-46 ventral $0.040$ ASD 1.975
Ventral Area 24d Area posterior 9-46 ventral $0.041$ ASD 1.975
Area 2 Area posterior 9-46 ventral $0.040$ ASD 1.963
ProStriate Area Area posterior 9-46 ventral $0.041$ ASD 1.960
RetroInsular Cortex Area posterior 9-46 ventral $0.041$ ASD 1.960
Area 9 anterior Area posterior 9-46 ventral $0.041$ ASD 1.954
VentroMedial Visual Area 3 Area posterior 9-46 ventral $0.042$ ASD 1.953
ParaHippocampal Area 2 Area posterior 9-46 ventral $0.042$ ASD 1.947
Area anterior 9-46 ventral Area posterior 9-46 ventral $0.043$ ASD 1.941
Medial Belt Complex Area posterior 9-46 ventral $0.043$ ASD 1.940
Insular Granular Complex Area posterior 9-46 ventral $0.043$ ASD 1.936
Eighth Visual Area Area posterior 9-46 ventral $0.044$ ASD 1.933
Superior Frontal Language Area Area posterior 9-46 ventral $0.044$ ASD 1.931
Area TemporoParietoOccipital Junction 3 Area posterior 9-46 ventral $0.044$ ASD 1.921
Area 43 Area posterior 9-46 ventral $0.044$ ASD 1.920
Posterior InferoTemporal complex Area posterior 9-46 ventral $0.045$ ASD 1.915
Area 7PC Area posterior 9-46 ventral $0.045$ ASD 1.913
Area STSv posterior Area posterior 9-46 ventral $0.044$ ASD 1.912
Area V3B Area posterior 9-46 ventral $0.046$ ASD 1.908
Area Lateral IntraParietal dorsal Area posterior 9-46 ventral $0.045$ ASD 1.905
Area Lateral Occipital 1 Area posterior 9-46 ventral $0.046$ ASD 1.904
Area Posterior Insular 1 Area posterior 9-46 ventral $0.046$ ASD 1.900
Dorsal Transitional Visual Area Area posterior 9-46 ventral $0.046$ ASD 1.900
Fusiform Face Complex Area posterior 9-46 ventral $0.047$ ASD 1.890
Second Visual Area Area posterior 9-46 ventral $0.048$ ASD 1.887
PeriSylvian Language Area Area posterior 9-46 ventral $0.047$ ASD 1.887
Area s32 Area posterior 9-46 ventral $0.047$ ASD 1.885
Area 46 Area posterior 9-46 ventral $0.047$ ASD 1.884
Hippocampus Area posterior 9-46 ventral $0.048$ ASD 1.880
VentroMedial Visual Area 2 Area posterior 9-46 ventral $0.048$ ASD 1.879
Area 8Av Area posterior 9-46 ventral $0.048$ ASD 1.877
Area OP4-PV Area posterior 9-46 ventral $0.048$ ASD 1.877
Posterior Insular Area 2 Area posterior 9-46 ventral $0.048$ ASD 1.876
Area TemporoParietoOccipital Junction 2 Area posterior 9-46 ventral $0.047$ ASD 1.876
Area 55b Area posterior 9-46 ventral $0.048$ ASD 1.876
Area 23d Area posterior 9-46 ventral $0.048$ ASD 1.874
PreSubiculum Area posterior 9-46 ventral $0.048$ ASD 1.871
Anterior IntraParietal Area Area posterior 9-46 ventral $0.048$ ASD 1.870
Area PFcm Area posterior 9-46 ventral $0.049$ ASD 1.869
Ventral Area 6 Area posterior 9-46 ventral $0.049$ ASD 1.867
Area 33 prime Area posterior 9-46 ventral $0.049$ ASD 1.866
Area STSd posterior Area posterior 9-46 ventral $0.048$ ASD 1.865
Parieto-Occipital Sulcus Area 1 Area posterior 9-46 ventral $0.049$ ASD 1.862
Seventh Visual Area Area posterior 9-46 ventral $0.050$ ASD 1.860
Area OP1-SII Area posterior 9-46 ventral $0.049$ ASD 1.860
Area STSd posterior Area Lateral IntraParietal dorsal $0.049$ ASD 1.838

### v2

@@ Appendix F Interpretability
Region 1 Region 2 Disorder/HC Corrected p-value Cohen’s d
Area 2 L Area 11l R HC 0.002 2.232
Area PFcm L Area 11l R HC 0.003 2.191
RetroInsular Cortex L Area 11l R HC 0.003 2.162
Ventral Area 24d L Area 11l R HC 0.003 2.162
Area IntraParietal 2 L Area 11l R HC 0.003 2.149
Area Posterior 24 prime R Area 11l R HC 0.003 2.147
Area Posterior 24 prime L Area 11l R HC 0.003 2.147
Ventral Area 24d R Area 11l R HC 0.003 2.145
Area 1 L Area 11l R HC 0.003 2.141
Area 44 R Area 11l R HC 0.003 2.14
Primary Auditory Cortex L Area 11l R HC 0.003 2.127
Area dorsal 32 R Area 11l R HC 0.004 2.11
Area Frontal Opercular 5 L Area 11l R HC 0.004 2.106
Area anterior 32 prime L Area 11l R HC 0.004 2.104
Parieto-Occipital Sulcus Area 2 R Area 11l R HC 0.004 2.101
Anterior IntraParietal Area L Area 11l R HC 0.004 2.101
Anterior 24 prime R Area 11l R HC 0.004 2.099
Area 43 R Area 11l R HC 0.004 2.097
Orbital Frontal Complex R Area 11l R HC 0.004 2.091
Area OP4-PV R Area 11l R HC 0.004 2.09
Area 23d L Area 11l R HC 0.004 2.085
Frontal Opercular Area 3 L Area 11l R HC 0.004 2.083
Area PGi R Area 11l R HC 0.004 2.082
Dorsal Area 24d L Area 11l R HC 0.004 2.082
Primary Sensory Cortex L Area 11l R HC 0.004 2.08
Second Visual Area R Sixth Visual Area L ASD 0.004 2.075
Area PFt L Area 11l R HC 0.004 2.075
Area 7PC L Area 11l R HC 0.004 2.073
Area anterior 32 prime R Area 11l R HC 0.004 2.07
Area 33 prime L Area 11l R HC 0.004 2.067
Area OP2-3-VS L Area 11l R HC 0.004 2.066
Area posterior 10p R Area 11l R HC 0.004 2.064
Dorsal Transitional Visual Area R Area 11l R HC 0.004 2.064
Area ventral 23 a+b L Area 11l R HC 0.004 2.055
Area 23d R Area 11l R HC 0.004 2.054
Area 7m R Area 11l R HC 0.004 2.053
Area 10v R Area 11l R HC 0.004 2.05
Lateral Area 7A L Area 11l R HC 0.004 2.05
Area IntraParietal 1 L Area 11l R HC 0.004 2.049
Area Frontal Opercular 5 R Area 11l R HC 0.004 2.049
Anterior 24 prime L Area 11l R HC 0.004 2.049
Area posterior 9-46v L Area 11l R HC 0.004 2.046
Area p32 prime R Area 11l R HC 0.004 2.046
Area ventral 23 a+b R Area 11l R HC 0.004 2.045
Area posterior 24 L Area 11l R HC 0.004 2.045
Area dorsal 32 L Area 11l R HC 0.004 2.044
PeriSylvian Language Area L Area 11l R HC 0.004 2.044
Auditory 4 Complex R Area 11l R HC 0.004 2.044
Lateral Belt Complex R Area 11l R HC 0.004 2.043
Area 8BM R Area 11l R HC 0.004 2.043
Area PGs L Area 11l R HC 0.004 2.041
Area 10d R Area 11l R HC 0.004 2.04
Fourth Visual Area L Sixth Visual Area L ASD 0.005 2.039
Rostral Area 6 R Area 11l R HC 0.005 2.038
Area 33 prime R Area 11l R HC 0.005 2.035
Insular Granular Complex L Area 11l R HC 0.005 2.034
Ventral Area 6 R Area 11l R HC 0.005 2.034
Seventh Visual Area R Sixth Visual Area L ASD 0.005 2.032
Area dorsal 23 a+b L Area 11l R HC 0.005 2.027
Orbital Frontal Complex L Area 11l R HC 0.005 2.027
Area TE2 anterior L Area 11l R HC 0.005 2.027
Medial Belt Complex R Area 11l R HC 0.005 2.026
Area 45 R Area 11l R HC 0.005 2.025
Medial Area 7P R Area 11l R HC 0.005 2.023
Ventral Area 6 L Area 11l R HC 0.005 2.021
Dorsal Area 24d R Area 11l R HC 0.005 2.02
ProStriate Area L Area 11l R HC 0.005 2.02
Area OP1-SII L Area 11l R HC 0.005 2.02
Medial IntraParietal Area L Area 11l R HC 0.005 2.019
Area anterior 47r R Area 11l R HC 0.005 2.018
Superior 6-8 Transitional Area R Area 11l R HC 0.005 2.015
Area 8Ad R Area 11l R HC 0.005 2.012
Area IFSp L Area 11l R HC 0.005 2.012
Superior Temporal Visual Area L Area 11l R HC 0.005 2.011
Ventral IntraParietal Complex L Area 11l R HC 0.005 2.011
Area p32 prime L Area 11l R HC 0.005 2.01
PreSubiculum R Sixth Visual Area L ASD 0.005 2.007
Area 8B Lateral R Area 11l R HC 0.005 2.005
Area V6A L Area 11l R HC 0.005 2.004
Area IFJp L Area 11l R HC 0.005 2.004
Area Lateral IntraParietal ventral L Area 11l R HC 0.005 2.004
Area 8Av R Area 11l R HC 0.005 2.004
Area IFJa L Area 11l R HC 0.005 2.004
Parieto-Occipital Sulcus Area 2 L Area 11l R HC 0.005 2.003
Frontal Opercular Area 4 L Area 11l R HC 0.005 2.003
Area PF Opercular L Area 11l R HC 0.005 2.001
Area posterior 24 R Area 11l R HC 0.005 2.001
Area PFm Complex R Area 11l R HC 0.005 2.0
Area 5L R Area 11l R HC 0.005 2.0
Area 10v L Area 11l R HC 0.005 1.998
Area 5m ventral L Area 11l R HC 0.005 1.998
Anterior Ventral Insular Area L Area 11l R HC 0.005 1.996
Area posterior 9-46v R Area 11l R HC 0.005 1.994
Area PGp L Area 11l R HC 0.005 1.992
Frontal Opercular Area 2 L Area 11l R HC 0.005 1.991
Area PFm Complex L Area 11l R HC 0.005 1.99
Area Lateral IntraParietal dorsal L Area 11l R HC 0.005 1.988
Parieto-Occipital Sulcus Area 1 L Area 11l R HC 0.005 1.988
Area 31p ventral L Area 11l R HC 0.005 1.987
Area V6A R Area 11l R HC 0.005 1.987
@@
Frontal Opercular Area 3 L Polar 10p R HC 0.003 2.128
Frontal Opercular Area 2 L Polar 10p R HC 0.003 2.123
Area V3A R Polar 10p R HC 0.003 2.117
Area posterior 47r L Polar 10p R HC 0.004 2.109
IntraParietal Sulcus Area 1 L Polar 10p R HC 0.004 2.101
Area OP2-3-VS R Polar 10p R HC 0.004 2.1
Frontal Opercular Area 3 R Polar 10p R HC 0.004 2.096
Frontal Opercular Area 2 R Polar 10p R HC 0.004 2.095
Area 44 L Polar 10p R HC 0.004 2.093
Para-Insular Area R Polar 10p R HC 0.004 2.093
Eighth Visual Area R Polar 10p R HC 0.004 2.091
Ventral IntraParietal Complex R Polar 10p R HC 0.004 2.09
Area IFSp L Polar 10p R HC 0.004 2.087
Area anterior 9-46v L Polar 10p R HC 0.004 2.083
Area STGa R Polar 10p R HC 0.004 2.082
Medial IntraParietal Area L Polar 10p R HC 0.004 2.078
ParaHippocampal Area 1 R Polar 10p R HC 0.004 2.075
Rostral Area 6 R Polar 10p R HC 0.004 2.074
Area V3B L Polar 10p R HC 0.004 2.074
Pirform Cortex L Polar 10p R HC 0.004 2.074
Anterior 24 prime L Polar 10p R HC 0.004 2.073
Para-Insular Area L Polar 10p R HC 0.004 2.072
Area TF R Polar 10p R HC 0.004 2.071
Primary Auditory Cortex R Polar 10p R HC 0.004 2.071
Area PFt R Polar 10p R HC 0.004 2.069
Area V4t L Polar 10p R HC 0.004 2.069
Area OP1-SII R Polar 10p R HC 0.004 2.069
Area V3B R Polar 10p R HC 0.004 2.069
Pirform Cortex R Polar 10p R HC 0.004 2.068
Area IntraParietal 0 L Polar 10p R HC 0.004 2.067
Area Frontal Opercular 5 L Polar 10p R HC 0.004 2.064
Frontal Opercular Area 1 R Polar 10p R HC 0.004 2.064
Area 55b L Polar 10p R HC 0.004 2.064
Area OP2-3-VS L Polar 10p R HC 0.004 2.064
ProStriate Area R Polar 10p R HC 0.004 2.063
Area STSd posterior L Polar 10p R HC 0.004 2.062
Area IntraParietal 0 R Polar 10p R HC 0.004 2.062
Area 8B Lateral L Polar 10p R HC 0.004 2.061
Area STSd anterior L Polar 10p R HC 0.004 2.061
Lateral Belt Complex R Polar 10p R HC 0.004 2.061
Lateral Area 7P L Polar 10p R HC 0.004 2.061
IntraParietal Sulcus Area 1 R Polar 10p R HC 0.004 2.06
Area OP1-SII L Polar 10p R HC 0.004 2.06
Area 2 R Polar 10p R HC 0.004 2.059
Area IFJa R Polar 10p R HC 0.004 2.059
Ventral IntraParietal Complex L Polar 10p R HC 0.004 2.058
Area Posterior Insular 1 R Polar 10p R HC 0.004 2.058
Area 7PC R Polar 10p R HC 0.004 2.058
Area IFJp R Polar 10p R HC 0.004 2.057
Insular Granular Complex R Polar 10p R HC 0.004 2.057
Area IFSa L Polar 10p R HC 0.004 2.057
Dorsal Transitional Visual Area L Polar 10p R HC 0.004 2.056
Posterior Insular Area 2 R Polar 10p R HC 0.004 2.055
Area 33 prime R Polar 10p R HC 0.004 2.055
Frontal Opercular Area 4 L Polar 10p R HC 0.004 2.054
Area Lateral IntraParietal ventral R Polar 10p R HC 0.004 2.053
Area 8BM L Polar 10p R HC 0.004 2.053
Ventral Visual Complex R Polar 10p R HC 0.004 2.053
Middle Insular Area L Polar 10p R HC 0.004 2.053
Area 9 Middle R Polar 10p R HC 0.004 2.052
Insular Granular Complex L Polar 10p R HC 0.004 2.051
Sixth Visual Area R Polar 10p R HC 0.004 2.051
Lateral Area 7A R Polar 10p R HC 0.004 2.051
Ventral Visual Complex L Polar 10p R HC 0.004 2.05
Area TG dorsal R Polar 10p R HC 0.004 2.05
Frontal Opercular Area 4 R Polar 10p R HC 0.004 2.05
Anterior IntraParietal Area R Polar 10p R HC 0.004 2.05
Hippocampus R Polar 10p R HC 0.004 2.05
PreSubiculum L Polar 10p R HC 0.004 2.05
Area anterior 10p L Polar 10p R HC 0.004 2.048
Area IFJp L Polar 10p R HC 0.004 2.047
Seventh Visual Area R Polar 10p R HC 0.004 2.047
VentroMedial Visual Area 2 R Polar 10p R HC 0.004 2.046
Eighth Visual Area L Polar 10p R HC 0.004 2.045
Area TA2 R Polar 10p R HC 0.004 2.044
Fourth Visual Area R Polar 10p R HC 0.004 2.044
Area PFcm R Polar 10p R HC 0.004 2.044
Area PFm Complex R Polar 10p R HC 0.004 2.044
Area PFm Complex L Polar 10p R HC 0.004 2.044
Area TE1 anterior R Polar 10p R HC 0.004 2.043
Posterior Insular Area 2 L Polar 10p R HC 0.004 2.043
Area 8Ad L Polar 10p R HC 0.004 2.043
Area dorsal 32 R Polar 10p R HC 0.004 2.042
Area V4t R Polar 10p R HC 0.004 2.042
Frontal Eye Fields L Polar 10p R HC 0.004 2.041
Posterior InferoTemporal complex L Polar 10p R HC 0.004 2.041
Area Lateral IntraParietal dorsal R Polar 10p R HC 0.004 2.041
Area PF Complex R Polar 10p R HC 0.004 2.04
Area 9 Middle L Polar 10p R HC 0.004 2.04
Area STSd anterior R Polar 10p R HC 0.004 2.04
Area 43 L Polar 10p R HC 0.004 2.04
VentroMedial Visual Area 3 L Polar 10p R HC 0.004 2.04
Area 8Av L Polar 10p R HC 0.004 2.039
Area 6mp L Polar 10p R HC 0.004 2.039
Area TE2 anterior R Polar 10p R HC 0.004 2.039
Area 1 L Polar 10p R HC 0.004 2.037
Area V6A L Polar 10p R HC 0.005 2.037
ParaBelt Complex R Polar 10p R HC 0.005 2.037
Area 44 R Polar 10p R HC 0.005 2.036
VentroMedial Visual Area 3 R Polar 10p R HC 0.005 2.036